# Multi-modal atmospheric characterization of Pictoris b

### v1

@@ Acknowledgements.
- Nowak et al. (2020) Nowak, M., Lacour, S., Lagrange, A. M., et al. 2020, Astronomy and Astrophysics, 642, L2, aDS Bibcode: 2020A&A…642L…2N
@@
- Parker et al. (2024) Parker, L. T., Birkby, J. L., Landman, R., et al. 2024, Monthly Notices of the Royal Astronomical Society, 531, 2356, publisher: OUP ADS Bibcode: 2024MNRAS.531.2356P
- Petrus et al. (2020) Petrus, S., Bonnefoy, M., Chauvin, G., et al. 2020, Astronomy and Astrophysics, 633, A124, aDS Bibcode: 2020A&A…633A.124P
- Petrus et al. (2021) Petrus, S., Bonnefoy, M., Chauvin, G., et al. 2021, Astronomy and Astrophysics, 648, A59, aDS Bibcode: 2021A&A…648A..59P
- Petrus et al. (2023) Petrus, S., Chauvin, G., Bonnefoy, M., et al. 2023, Astronomy and Astrophysics, 670, L9, aDS Bibcode: 2023A&A…670L…9P
- Petrus et al. (2024) Petrus, S., Whiteford, N., Patapis, P., et al. 2024, The Astrophysical Journal, 966, L11, publisher: IOP ADS Bibcode: 2024ApJ…966L..11P
@@
- Quanz et al. (2010) Quanz, S. P., Meyer, M. R., Kenworthy, M. A., et al. 2010, The Astrophysical Journal, 722, L49, publisher: IOP ADS Bibcode: 2010ApJ…722L..49Q
@@
- Rotman et al. (2025) Rotman, Y., Welbanks, L., Line, M. R., et al. 2025, arXiv e-prints, arXiv:2503.21702
- Ruffio et al. (2019) Ruffio, J.-B., Macintosh, B., Konopacky, Q. M., et al. 2019, The Astronomical Journal, 158, 200, publisher: IOP ADS Bibcode: 2019AJ….158..200R
- Samland et al. (2017) Samland, M., Mollière, P., Bonnefoy, M., et al. 2017, Astronomy and Astrophysics, 603, A57, aDS Bibcode: 2017A&A…603A..57S
@@
- Skilling (2004) Skilling, J. 2004, 735, 395, conference Name: Bayesian Inference and Maximum Entropy Methods in Science and Engineering: 24th International Workshop on Bayesian Inference and Maximum Entropy Methods in Science and Engineering Publisher: AIP ADS Bibcode: 2004AIPC..735..395S
- Smith & Terrile (1984) Smith, B. A. & Terrile, R. J. 1984, Science, 226, 1421, aDS Bibcode: 1984Sci…226.1421S
- Snellen et al. (2014) Snellen, I., Brandl, B., de Kok, R., et al. 2014, The fast spin-rotation of a young extrasolar planet, publication Title: arXiv e-prints ADS Bibcode: 2014arXiv1404.7506S
- Snellen et al. (2014) Snellen, I. A. G., Brandl, B. R., de Kok, R. J., et al. 2014, Nature, 509, 63
@@
- Swastik et al. (2021) Swastik, C., Banyal, R. K., Narang, M., et al. 2021, The Astronomical Journal, 161, 114, publisher: IOP ADS Bibcode: 2021AJ….161..114S
- Tremblin et al. (2015) Tremblin, P., Amundsen, D. S., Mourier, P., et al. 2015, The Astrophysical Journal, 804, L17, publisher: IOP ADS Bibcode: 2015ApJ…804L..17T
@@
- Worthen et al. (2024) Worthen, K., Chen, C. H., Law, D. R., et al. 2024, The Astrophysical Journal, 964, 168, publisher: IOP ADS Bibcode: 2024ApJ…964..168W
@@
- Zhang et al. (2021) Zhang, Y., Snellen, I. A. G., Bohn, A. J., et al. 2021, Nature, 595, 370, aDS Bibcode: 2021Natur.595..370Z
- Zhang et al. (2024) Zhang, Y., Xuan, J. W., Mawet, D., et al. 2024, arXiv e-prints, arXiv:2407.20952
@@
- Zucker (2003) Zucker, S. 2003, Monthly Notices of the Royal Astronomical Society, 342, 1291, publisher: OUP ADS Bibcode: 2003MNRAS.342.1291Z
- Zwintz et al. (2019) Zwintz, K., Reese, D. R., Neiner, C., et al. 2019, Astronomy and Astrophysics, 627, A28, aDS Bibcode: 2019A&A…627A..28Z
@@ Appendix A Simulation of 13CO detection
This section summarizes the tests we performed using synthetic spectra to constrained both the value and detection limits of the isotopologue 13CO using GRAVITY. We generated two classes synthetic spectra using the Exo-REM grid with similar parameters as the ones predicted for Pic b (namely Chilcote et al. (2017)) both with and without 13CO and at the resolution, binning and wavelength coverage of GRAVITY. We then created our data samples divided by signal-to-noise: low (S/N ), medium (S/N , similar to the GRAVITY data) and high (S/N ). This noise was effectively applied to the data using this formula:
where is the original synthetic spectrum, a Gaussian distribution centered on 0 with a standard deviation of and final noised spectrum. This effectively scale the error bars with flux to obtain an overall constant S/N across the K-band.
@@
Obviously, this is an idealized case where the synthetic observation is only affected by Gaussian (photon) noise. To better mimic the effect of the atmospheric residuals seen on the GRAVITY spectrum, we modulated the synthetic flux with an observed transmission spectrum (T) taken from ESO/SkyCalc. This modulation was applied using the following formula:
@@
Appendix B GRAVITY analysis
Reference Figures of Sect. 4.
Appendix C Multi-modal analysis
Reference Figures and Tables of Sect. 5
| Parameter | |||||||
@@ Appendix C Multi-modal analysis
Appendix D Literature data on Pictoris b
This section is a summary of this existing literature on Pic b. Table 12 aim to be an exhaustive list of the physical parameter estimations coming from both atmospheric (forward modeling and retrieval) and evolutionary models.
| Reference | Data | Model | ||||||
@@ Appendix D Literature data on Pictoris b
| Houllé et al. (2024, subm.) | MATISSE, GRAVITY fixed | Exo-REM | ||||||

### v2

@@ Acknowledgements.
- Nowak et al. (2020) Nowak, M., Lacour, S., Lagrange, A. M., et al. 2020, A&A, 642, L2
@@
- Parker et al. (2024) Parker, L. T., Birkby, J. L., Landman, R., et al. 2024, MNRAS, 531, 2356
- Petrus et al. (2020) Petrus, S., Bonnefoy, M., Chauvin, G., et al. 2020, A&A, 633, A124
- Petrus et al. (2021) Petrus, S., Bonnefoy, M., Chauvin, G., et al. 2021, A&A, 648, A59
- Petrus et al. (2023) Petrus, S., Chauvin, G., Bonnefoy, M., et al. 2023, A&A, 670, L9
- Petrus et al. (2025) Petrus, S., Chauvin, G., Bonnefoy, M., et al. 2025, A&A, 701, A208
- Petrus et al. (2024) Petrus, S., Whiteford, N., Patapis, P., et al. 2024, ApJ, 966, L11
@@
- Quanz et al. (2010) Quanz, S. P., Meyer, M. R., Kenworthy, M. A., et al. 2010, ApJ, 722, L49
@@
- Rotman et al. (2025) Rotman, Y., Welbanks, L., Line, M. R., et al. 2025, ApJ, 989, 201
- Ruffio et al. (2019) Ruffio, J.-B., Macintosh, B., Konopacky, Q. M., et al. 2019, AJ, 158, 200
- Samland et al. (2017) Samland, M., Mollière, P., Bonnefoy, M., et al. 2017, A&A, 603, A57
@@
- Skilling (2004) Skilling, J. 2004, in American Institute of Physics Conference Series, Vol. 735, Bayesian Inference and Maximum Entropy Methods in Science and Engineering: 24th International Workshop on Bayesian Inference and Maximum Entropy Methods in Science and Engineering, ed. R. Fischer, R. Preuss, & U. V. Toussaint (AIP), 395–405
- Smith & Terrile (1984) Smith, B. A. & Terrile, R. J. 1984, Science, 226, 1421
- Snellen et al. (2014a) Snellen, I., Brandl, B., de Kok, R., et al. 2014a, arXiv e-prints, arXiv:1404.7506
- Snellen et al. (2014b) Snellen, I. A. G., Brandl, B. R., de Kok, R. J., et al. 2014b, Nature, 509, 63
@@
- Swastik et al. (2021) Swastik, C., Banyal, R. K., Narang, M., et al. 2021, AJ, 161, 114
- Tremblin et al. (2015) Tremblin, P., Amundsen, D. S., Mourier, P., et al. 2015, ApJ, 804, L17
@@
- Worthen et al. (2024) Worthen, K., Chen, C. H., Law, D. R., et al. 2024, ApJ, 964, 168
@@
- Zhang et al. (2021) Zhang, Y., Snellen, I. A. G., Bohn, A. J., et al. 2021, Nature, 595, 370
- Zhang et al. (2024) Zhang, Y., Xuan, J. W., Mawet, D., et al. 2024, AJ, 168, 131
@@
- Zucker (2003) Zucker, S. 2003, MNRAS, 342, 1291
- Zwintz et al. (2019) Zwintz, K., Reese, D. R., Neiner, C., et al. 2019, A&A, 627, A28
@@ Appendix A Simulation of 13CO detection
This section summarizes the tests we performed using synthetic spectra to constrained both the value and detection limits of the isotopologue 13CO using GRAVITY. We generated two classes synthetic spectra using the Exo-REM grid with similar parameters as the ones predicted for Pic b (namely Chilcote et al. (2017)) both with and without 13CO and at the resolution, binning and wavelength coverage of GRAVITY. We then created our data samples divided by signal-to-noise: low (S/N ), medium (S/N , similar to the GRAVITY data) and high (S/N ). This noise was effectively applied to the data using this formula,
where is the original synthetic spectrum, a Gaussian distribution centered on 0 with a standard deviation of and final noised spectrum. This effectively scale the error bars with flux to obtain an overall constant S/N across the K band.
@@
Obviously, this is an idealized case where the synthetic observation is only affected by Gaussian (photon) noise. To better mimic the effect of the atmospheric residuals seen on the GRAVITY spectrum, we modulated the synthetic flux with an observed transmission spectrum (T) taken from ESO/SkyCalc. This modulation was applied using the following formula,
@@
Appendix B Figures and tables
| Parameter | log(g) | [M/H] | C/O | f | log(12C/13C) | ||
| (K) | (dex) | ||||||
@@ Appendix B Figures and tables
| Reference | Data | Model | log(g) | [M/H] | C/O | R | M | |
| (K) | (dex) | () | () | |||||
@@
| Houllé et al. (2025) | MATISSE, GRAVITY fixed | Exo-REM | ||||||